### v1

# Symmetry restoration for TDiff scalar fields

Abstract
We explore the idea of restoring the full diffeomorphism (Diff) invariance in theories with only transverse diffeomorphisms (TDiff) by the introduction of additional fields. In particular, we consider in detail the case of a TDiff invariant scalar field and how Diff symmetry can be restored by introducing an additional vector field. We reobtain the corresponding dynamics and energy-momentum tensor from the covariantized action and analyze the potential and kinetic domination regimes. For the former, the theory describes a cosmological constant-type behaviour, while for the latter we show that the theory can describe an adiabatic perfect fluid whose equation of state and speed of sound is obtained in a straightforward way.
preprint: IPARCOS-UCM-24-013
I Introduction
Our current best description of gravitational phenomena is, and has been for over a century, the theory of General Relativiy (GR). It is not a theory without weaknesses, however, and although it performs remarkably well in numerous tests on solar system scales, there is reason to believe that it is not the end of the story. For one, it is not a theory which serves to describe gravity in its most extreme regimes, where one expects quantum effects to gain importance, but it is also possible to find problems even while remaining classical. Indeed, the various tensions between theory and observations in cosmology are another hint at the possibility that the theory breaks down at such scales. Sparked by considerations of the sort, together with other theoretical issues such as the cosmological constant problem, modified theories of gravity have been a central object of study in this regard (see e.g. [1] for a review).
In particular, it proves worthwhile to reconsider the fundamental symmetries involved in GR, namely the diffeomorphism (Diff) invariance of the theory. This amounts to the assertion that the physical equations remain invariant under general coordinate transformations. The study of situations where such a symmetry is broken can in fact be traced back to Einstein himself in 1919 with the introduction of Unimodular Gravity [2], where the metric determinant is reduced to be a non-dynamical field fixed to the value $g=1$. Unimodular Gravity is perhaps the most well-known example of a theory with broken Diff invariance, being the symmetry group in that case the union of transverse diffeomorphisms (TDiff) and Weyl rescalings (together dubbed WTDiff, see e.g. [3] for a review). The equations of motion of the theory are the trace-free Einstein equations (see e.g. [4] for a comprehensive introduction), in which any cosmological constant-type contribution does not gravitate, thus providing an elegant solution to the cosmological constant problem.
In more recent years, interest has grown in theories which present TDiff invariance. Simply put, transverse diffeomorphisms are general coordinate transformations in which the Jacobian determinant is required to be $J=1$. Infinitesimally, if we consider the coordinate transformation $x^{\mu}\rightarrow\hat{x}^{\mu}=x^{\mu}+\xi^{\mu}(x)$ generated by a vector field $\xi^{\mu}(x)$, then what we do is require the condition $\partial_{\mu}\xi^{\mu}=0$ (see [5] for a concise introduction to transverse diffeomorphisms). The fact that the Jacobian determinant equals unity means that objects which were tensor densities under Diff become actual tensors under TDiff. In particular, the metric determinant becomes a TDiff scalar field, and this has interesting implications. Indeed, on the one hand, the metric determinant becomes a scalar field to which one may endow dynamics. On the other hand, the invariant volume element we find in an action integral is no longer fixed by the symmetry to be $d\text{vol}=\sqrt{g}\,d^{4}x$, but can actually take on the more general form $d\text{vol}=f(g)\,d^{4}x$, with $f(g)$ an arbitrary function of the metric determinant, and this opens up an enormous range of possibilities for novel couplings. Field theories in which the gravitational sector is TDiff invariant are studied in references [6, 7, 8, 9, 10], where cosmological implications are also discussed. One can also study the consequences of breaking the symmetries in the matter sector. TDiff invariant theories with a scalar field were recently considered in references [11, 5, 12, 13]. Reference [5] considers general scalar field TDiff theories in cosmological contexts, reference [12] performs a general study for a scalar field without assuming any background geometry, and reference [13] provides a unified description for the dark sector using a particular theory, comparing the results with the latest cosmological observations and datasets.
Now, it is not an uncommon situation in physics to find several equivalent descriptions, or reformulations, of the same theory. Examples include, but are by no means limited to, the well-known equivalence between the Palatini and the metric approach to GR, the scalar-tensor perspective of $f(R)$-gravity (see, for example, [1]), the correspondence between the field equations of non linear Ricci-based metric-affine theories of gravity coupled to scalar matter and GR coupled to a different scalar field Lagrangian [14]. Moreover, ever since the pioneering work of Stueckelberg [15] (see e.g. [16] for a review), it is well known how a theory may recover (or reveal) its broken gauge symmetry via the introduction of additional fields. With these two ideas in mind, we present in this work an alternative formalism for the treatment of a TDiff theory for a scalar field based on previous work done by Henneaux and Teitelboim [17] within the context of Unimodular Gravity (see also Kuchar [18] for an alternative approach). Following that spirit, we shall in this work show how a TDiff invariant theory may be equivalently described as a Diff invariant theory with an additional field. We will then work on the particular case of a scalar field theory from both points of view. This idea of finding an equivalent description with symmetry restoration has been also applied on the different gravitational framework of massive gravity [19].
The paper is organized as follows. First of all, in section II we review the TDiff approach for the scalar field, and present the main results. In section III we show how it is possible to reformulate a TDiff invariant theory in a way which recovers Diff invariance via the introduction of an additional field. Section IV is then devoted to the reformulation of our scalar field theory in a covariantized manner. Finally, in section V we present the main conclusions of the work and discuss further outlook. In Appendix A we have included some calculations which are not needed to follow the thread of the main discussion, but which the reader might find useful.
As a final note, we remark that our conventions in this work include the usage of units in which $\hbar=c=1$, metric signature (+,,,)(+,-,-,-)( + , - , - , - ), and the notation $g=\absolutevalue{\det(g_{\mu\nu})}$ |.
II The TDiff approach
The breaking of Diff invariance down to TDiff in the matter action was recently studied in general backgrounds in reference [12]. This work considers a scalar field coupled to gravity via arbitrary functions of the metric determinant, with total action
$S=S_{\text{EH}}+S_{m}\,,$ (1)
where
$S_{\text{EH}}=-\frac{1}{16\pi G}\int d^{4}x\,\sqrt{g}\,R\,,$ (2)
and
$S_{m}=\int d^{4}x\,\left\{\frac{f_{k}(g)}{2}g^{\mu\nu}\partial_{\mu}\psi\partial_{\nu}\psi-f_{v}(g)V(\psi)\right\}\,.$ (3)
This action is indeed seen to be invariant only under transverse diffeomorphisms due to the arbitrary functions of (the absolute value of) the metric determinant $f_{k}(g)$ and $f_{v}(g)$.
Let us summarize in this section the main results obtained in reference [12] following the TDiff approach. The equation of motion (EoM) for the scalar field is
$\partial_{\mu}\big{(}f_{k}(g)\partial^{\mu}\psi\big{)}+f_{v}(g)V^{\prime}(\psi)=0\,,$ (4)
where in general a prime denotes derivative with respect to its argument. On the other hand, the EoM for the gravitational field are the usual Einstein equations
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi G\,T_{\mu\nu}\,,$ (5)
where the Energy-Momentum Tensor (EMT) for the scalar field is found from its definition
$T_{\mu\nu}=\frac{2}{\sqrt{g}}\frac{\delta S_{m}}{\delta g^{\mu\nu}}\,,$ (6)
and reads:
$\begin{split}T_{\mu\nu}&=\frac{2}{\sqrt{g}}\left\{\frac{1}{2}f_{k}(g)\partial_{\mu}\psi\partial_{\nu}\psi\right.+\\[5.0pt] &\quad+\left.g\left[f_{v}^{\prime}(g)V(\psi)-\frac{1}{2}f^{\prime}_{k}(g)(\partial\psi)^{2}\right]g_{\mu\nu}\right\}\,,\end{split}$ (7)
where we are denoting
$(\partial\psi)^{2}\equiv\partial_{\alpha}\psi\partial^{\alpha}\psi\,.$ (8)
Under the assumption of the field derivative $\partial_{\mu}\psi$ being a timelike vector, it is possible to rewrite the EMT in perfect fluid form. Indeed, defining a unit timelike vector field $u^{\mu}$ through
$u^{\mu}=\frac{\partial^{\mu}\psi}{\sqrt{(\partial\psi)^{2}}}\equiv\frac{\partial^{\mu}\psi}{N}\,,$ (9)
where we are denoting the normalization as
$N\equiv\sqrt{(\partial\psi)^{2}}\,,$ (10)
together with an energy density
$\rho=\frac{2}{\sqrt{g}}\left\{\frac{1}{2}f_{k}(\partial\psi)^{2}+g\left[f_{v}^{\prime}V-\frac{1}{2}f^{\prime}_{k}(\partial\psi)^{2}\right]\right\}$ (11)
and a pressure
$p=-\frac{2g}{\sqrt{g}}\left[f_{v}^{\prime}V-\frac{1}{2}f^{\prime}_{k}(\partial\psi)^{2}\right]\,,$ (12)
one can rewrite the EMT (7) as
$T_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}-pg_{\mu\nu}\,.$ (13)
Since our intention is reformulating some previously done study, we shall in this paper also work with the timelike vector assumption whenever we wish to reexpress the analysis as that of a perfect fluid.
One of the main points of study in the TDiff approach is the conservation of the EMT. Indeed, the conservation of this quantity is an automatic consequence of the Noether theorem for a theory with symmetry under diffeomorphisms, but it does not follow trivially when we have less symmetry. Instead, one argues that the conservation of the EMT on the solutions to the equations of motion is a consistency requirement of the theory, since the Einstein equations (5) still hold and the Einstein tensor is divergenceless:
$\nabla_{\mu}G^{\mu\nu}=0\,\implies\,\nabla_{\mu}T^{\mu\nu}=0\,.$ (14)
Within the TDiff approach, this consistency condition allows one to obtain a certain (physical) constraint on the metric.
We shall now focus on two limiting cases of interest. These are the potential domination regime and the kinetic domination regime, which we review in the following.
II.1 Potential domination in the TDiff approach
Everything is rather simple in the potential regime, which amounts to ignoring the kinetic contribution in the action. When we do so, the EoM (4) for $\psi$ becomes $f_{v}V^{\prime}=0$, which (for a nonvanishing coupling function $f_{v}$) tells us that the field takes on the value $\psi=\psi_{0}$ which is the extremum of the potential: $V(\psi)=V(\psi_{0})=$ const. The EMT (7) simplifies down to
$T_{\mu\nu}=2Vf_{v}^{\prime}\sqrt{g}\,g_{\mu\nu}\,,$ (15)
and its conservation becomes
$0=\nabla_{\mu}T^{\mu\nu}=2Vg^{\mu\nu}\nabla_{\mu}\left(f_{v}^{\prime}\sqrt{g}\right)=2Vg^{\mu\nu}\partial_{\mu}\left(f_{v}^{\prime}\sqrt{g}\right)\,,$ (16)
where in the first equality we have pulled (covariantly) constant terms out of the covariant derivative and in the second equality we have recognized that the product $f_{v}^{\prime}\sqrt{g}$ is a TDiff scalar so that we may use a partial derivative. Since relation (16) must be met for all metrics, it follows that
$\partial_{\mu}\left(f_{v}^{\prime}\sqrt{g}\right)=\frac{1}{\sqrt{g}}\left(\frac{1}{2}f^{\prime}_{v}+gf_{v}^{\prime\prime}\right)\partial_{\mu}g=0\,.$ (17)
The above relation is satisfied whenever the coupling function takes the form
$f_{v}(g)=A\sqrt{g}+B\,,$ (18)
with $A$ and $B$ constants of integration, but if we wish to leave the coupling function arbitrary (which in principle we do), then it must be the case that
$\partial_{\mu}g=0\,\implies\,g=\text{constant}\,.$ (19)
This is the constraint on the metric which we obtain in the potential domination regime: the determinant must be constant. It is interesting to note that, being the determinant a constant quantity, any given function of the determinant will also assume a constant value, for instance the function $f_{v}^{\prime}(g)$. This is relevant because, looking back at the EMT (15), it may be written as
$T_{\mu\nu}\equiv\lambda\,g_{\mu\nu}\,,$ (20)
with
$\lambda\equiv 2Vf_{v}^{\prime}\sqrt{g}=\text{constant}\,,$ (21)
and we have the behaviour of a cosmological constant.
II.2 Kinetic domination in the TDiff approach
The study of the kinetic domination regime is more involved. The EoM for the field in this regime becomes $\partial_{\mu}\left(f_{k}\partial^{\mu}\psi\right)=0$, and its solution must satisfy
$(\partial\psi)^{2}=C_{\psi}(x)\left(\frac{\sqrt{g}}{\delta Vf_{k}}\right)^{2}\,,$ (22)
with $C_{\psi}(x)$ a function such that $u^{\mu}\partial_{\mu}C_{\psi}(x)=0$, and where we find the cross-sectional volume of the congruence $\delta V$, related to the expansion by [20]
$\nabla_{\mu}u^{\mu}=u^{\mu}\partial_{\mu}\ln\delta V\,.$ (23)
[…]
$\dot{\rho}+(\rho+p)\nabla_{\mu}u^{\mu}=0\,,$ (24a)
$(\rho+p)\dot{u}^{\mu}-\left(g^{\mu\nu}-u^{\mu}u^{\nu}\right)\nabla_{\nu}p=0\,,$ (24b)
where we use the dot notation ˙$\dot{\,\,}\equiv u^{\mu}\nabla_{\mu}$. In the kinetic regime, the perfect fluid quantities take on the form
$\rho$ $=\frac{(\partial\psi)^{2}}{\sqrt{g}}\left(f_{k}-gf^{\prime}_{k}\right)\,,$ (25a)
$p$ $=\frac{(\partial\psi)^{2}}{\sqrt{g}}\,gf^{\prime}_{k}\,,$ (25b)
and the equation of state (EoS) reads
$w=\frac{p}{\rho}=\frac{gf^{\prime}_{k}}{f_{k}-gf^{\prime}_{k}}\equiv\frac{F}{1-F}\,,$ (26)
where we are defining
$F\equiv\frac{gf^{\prime}_{k}}{f_{k}}\,.$ (27)
It is interesting to remark that the EoS (26) is a function of only the metric determinant, $w=w(g)$.
Studying the longitudinal projection (24a) on the solution to the EoM (22) yields the following relation:
$(2F-1)\frac{g}{f_{k}}=C_{g}(x)\delta V^{2}\,,$ (28)
with $C_{g}(x)$ a function which must satisfy
$u^{\mu}\partial_{\mu}C_{g}(x)=0\,.$ (29)
Equation (28) is one of the main results in the TDiff approach, and indeed shows how the study of EMT conservation is not a trivial matter but rather imposes a constraint on the metric.
Probing the transverse projection (24b) for further information ends up revealing that the two functions $C_{\psi}(x)$ and $C_{g}(x)$ are inversely proportional, i.e.
$C_{\psi}C_{g}=\text{constant}\equiv c_{\rho}\,.$ (30)
The reason for naming the constant as $c_{\rho}$ is because it is possible to find the following nice expression for the energy density in the kinetic regime:
$\rho=\frac{c_{\rho}}{(w-1)\sqrt{g}}\,.$ (31)
Since the EoS is only a function of the metric determinant, it follows that both the energy density and the pressure are functions of only the metric determinant, and this dependence on a single quantity reveals that we are dealing with an adiabatic fluid. This is another of the main results in the TDiff approach, and the speed of sound for the possible adiabatic perturbations (defined through $\delta p=c_{s}^{2}\,\delta\rho$) is found to take the form:
$c_{s}^{2}=-\frac{gf_{k}(f^{\prime}_{k}+2gf_{k}^{\prime\prime})}{f_{k}^{2}+(2gf_{k}^{\prime})^{2}-gf_{k}(5f_{k}^{\prime}+2gf_{k}^{\prime\prime})}\,.$ (32)
Thus concludes our summary of the general framework resulting from the consideration of a TDiff scalar field. For interesting consequences and phenonomenology of this theory we refer the reader to references [5, 12, 13].
III Covariantized action
Now we note that one can rewrite an action with broken Diff invariance in a Diff invariant way via the introduction of additional fields, similar in spirit to the Stueckelberg procedure in gauge theories [15]. We reproduce in this section the procedure outlined in reference [5] for a scalar field and in reference [12] for general matter fields $\Psi$. Let us consider the following TDiff invariant action:
$S_{\text{TDiff}}[g_{\mu\nu},\Psi]=\int d^{4}x\,\sum_{i}f_{i}(g)\,\mathcal{L}_{i}\left(g_{\mu\nu},\Psi,\partial_{\mu}\Psi\right)$ (33)
(where the index $i$ refers to each of the possible terms we may find in the action), with each $\mathcal{L}_{i}$ a Diff scalar and $f_{i}$ arbitrary coupling functions of (the absolute value of) the metric determinant. In order to perform the rewrite, we first of all need to introduce a Diff scalar density $\bar{\mu}$ which transforms as $\sqrt{g}$ under general coordinate transformations. Following [5], let us define the density to be
$\bar{\mu}=\partial_{\mu}\left(\sqrt{g}\,T^{\mu}\right)\,,$ (34)
with $T^{\mu}$ a Diff vector field. The quantity $\bar{\mu}$ indeed transforms as $\sqrt{g}$, since
$\frac{\bar{\mu}}{\sqrt{g}}=\frac{1}{\sqrt{g}}\partial_{\mu}\left(\sqrt{g}\,T^{\mu}\right)=\nabla_{\mu}T^{\mu}\equiv Y\,,$ (35)
where we have denoted the scalar $Y\equiv\nabla_{\mu}T^{\mu}$. It is interesting to remark that in the general Stueckelberg procedure the idea is to take the gauge functions and promote them to fields. Since 4-dimensional diffeomorphisms are generated by 4 functions, it is natural to introduce a vector field $T^{\mu}$. This being said, however, going from TDiff back to Diff actually only involves removing one condition (recall $J=1$, or $\partial_{\mu}\xi^{\mu}=0$), and thus it makes sense that the newly introduced vector field ends up appearing through the scalar combination $Y=\nabla_{\mu}T^{\mu}$.
Consider now the following Diff invariant action
$S_{\text{Diff}}[g_{\mu\nu},\Psi,T^{\mu}]=\int d^{4}x\,\sum_{i}\sqrt{g}\,H_{i}\left(Y\right)\mathcal{L}_{i}\left(g_{\mu\nu},\Psi,\partial_{\mu}\Psi\right)\,,$ (36)
which, at a first glance, seems to be describing a similar theory to that of action (33), but with an additional vector field $T^{\mu}$ field coming into play via the arbitrary functions $H_{i}(Y)=H_{i}(\nabla_{\mu}T^{\mu})$. For convenience, we shall express the arbitrary functions $H_{i}(Y)$ as
$H_{i}(Y)\equiv Yf_{i}(Y^{-2})\,,$ (37)
with arbitrary $f_{i}$, so that the action is rewritten as
$\begin{split}S_{\text{Diff}}&[g_{\mu\nu},\Psi,T^{\mu}]=\\[5.0pt] &=\int d^{4}x\,\sum_{i}\sqrt{g}\,Yf_{i}(Y^{-2})\mathcal{L}_{i}\left(g_{\mu\nu},\Psi,\partial_{\mu}\Psi\right)\,.\end{split}$ (38)
We are now in a position to make this Diff action completely equivalent to the TDiff action (33) above by going to a coordinate system where $\bar{\mu}=1$, i.e. $Y=1/\sqrt{g}$. Indeed, in such a coordinate system we have that
$\sqrt{g}\,H_{i}(Y)\bigg{|}_{\bar{\mu}=1}=\bar{\mu}\,f_{i}\left(g/\bar{\mu}^{2}\right)\bigg{|}_{\bar{\mu}=1}=f_{i}(g)\,,$ (39)
and so it immediately follows that
$S_{\text{Diff}}[g_{\mu\nu},\Psi,T^{\mu}]\bigg{|}_{\bar{\mu}=1}=S_{\text{TDiff}}[g_{\mu\nu},\Psi]\,.$ (40)
In this manner, we see how one may alternatively work within the TDiff approach or the covariantized approach since they are equivalent, as it is explicitly shown when choosing the gauge $\bar{\mu}=1$ (equivalently, $Y=1/\sqrt{g}$) in the Diff invariant action. Thus, the invariance under full diffeomorphisms has been restored due to the introduction of a Diff vector field $T^{\mu}$.
IV The covariantized approach
Let us take the Diff invariant action [5]
$S_{\text{Diff}}[g_{\mu\nu},\psi,T^{\mu}]=S_{\text{EH}}+\int d^{4}x\,\sqrt{g}\left[H_{k}(Y)X-H_{v}(Y)V\right]$ (41)
where, for simplicity, we are denoting the kinetic term by
$X\equiv\frac{1}{2}g^{\mu\nu}\partial_{\mu}\psi\partial_{\nu}\psi=\frac{1}{2}(\partial\psi)^{2}\,.$ (42)
If we set $\bar{\mu}=1$ and recognize the functions $f_{k}(g)$ and $f_{v}(g)$, we immediately realize that it has become precisely action (1) for a TDiff scalar field. Let us keep working now in the covariantized approach, and see what we obtain. Variations of action (41) with respect to the scalar field $\psi$ yield the following EoM
$\nabla_{\nu}\left[H_{k}(Y)g^{\mu\nu}\nabla_{\mu}\psi\right]+H_{v}(Y)V^{\prime}(\psi)=0\,.$ (43)
Recalling that the covariant divergence of a vector $V^{\alpha}$ may be expressed as
$\nabla_{\alpha}V^{\alpha}=\frac{1}{\sqrt{g}}\partial_{\alpha}\left(\sqrt{g}\,V^{\alpha}\right)\,,$ (44)
and using equation (39), one can see that in the $\bar{\mu}=1$ gauge this equation reduces to
$\frac{1}{\sqrt{g}}\partial_{\nu}[f_{k}(g)\partial^{\mu}\psi]+\frac{f_{v}(g)}{\sqrt{g}}V^{\prime}(\psi)=0\,,$ (45)
which is equivalent to the EoM (4) we found for $\psi$ in the TDiff approach.
On the other hand, variations of action (41) with respect to the vector field that has restored the Diff invariance, that is $T^{\mu}$, provide the following EoM:
$\partial_{\nu}\left[H_{k}^{\prime}(Y)X-H_{v}^{\prime}(Y)V\right]=0\,.$ (46)
Hence, there is a conserved quantity as a result of the derivative dependence of the field $T^{\mu}$ on action (41). One could think that these equations are new with respect to the TDiff formalism. However, it should be noted that in said formalism one has the equations coming from the conservation of the EMT that is no longer trivially satisfied in that framework, which are equations (14). So, as we will show, these equations must be equivalent.
Finally, variations with respect to the metric tensor $g_{\mu\nu}$ provide the Einstein equations
$G_{\mu\nu}=8\pi G\,T_{\mu\nu}\,,$ (47)
where the total EMT is found using definition (6) and takes the form
$\begin{split}T_{\mu\nu}&=H_{k}(Y)\partial_{\mu}\psi\partial_{\nu}\psi-\left[H_{k}(Y)X-H_{v}(Y)V\right]g_{\mu\nu}\\[5.0pt] &\qquad+Y\left[H_{k}^{\prime}(Y)X-H_{v}^{\prime}(Y)V\right]g_{\mu\nu}\end{split}$ (48)
(see Appendix A for the calculation of this quantity). It should be noted that the conservation of the EMT (48) has to be trivially satisfied when considering the EoM of the fields (43) and (46), due to the invariance of action (41) under general diffeomorphisms. Indeed, it can be explicitly checked that this is the case, and we refer the reader to Appendix A for the calculation.
On the other hand, noting the definition of the $H(Y)$ functions in equation (37), one can obtain the relation
$H^{\prime}(Y)=f\left(Y^{-2}\right)-2Y^{-2}\frac{df}{d\left(Y^{-2}\right)}\,,$ (49)
so that
$H^{\prime}(Y)\bigg{|}_{\bar{\mu}=1}=f(g)-2gf^{\prime}(g)\,.$ (50)
Taking into account this result, together with equation (39), it is immediate to verify that the covariantized EMT (48) reduces to the TDiff EMT (7) in the $\bar{\mu}=1$ frame.
This covariantized formalism has been applied to the FLRW model in reference [5], recovering the information provided by the constraint on the metric in the TDiff case. In the next section we will show that this is possible in general.
IV.1 Potential domination in the covariantized approach
We begin with the potential domination regime. Neglecting the kinetic parts, the EoM (43) for $\psi$ becomes
$H_{v}(Y)V^{\prime}(\psi)=0\,,$ (51)
which means that either $H_{v}(Y)=0$ (we do not consider this possibility) or that the field takes on the constant value $\psi(x)=\psi_{0}$ such that the potential reaches an extremum $V(\psi)=V(\psi_{0})=$ constant (exactly as we concluded in the TDiff approach). On the other hand, the EoM (46) for $T^{\mu}$ in the potential regime become
$\partial_{\mu}\left[H_{v}^{\prime}(Y)V\right]=0\,\Rightarrow\,\partial_{\mu}H_{v}^{\prime}(Y)=H_{v}^{\prime\prime}(Y)\,\partial_{\mu}Y=0\,,$ (52)
where we have already used that $V=$ constant to pull it out of the derivative. We now have a couple of options in order for this equation to be satisfied, namely that $H_{v}^{\prime\prime}(Y)=0$ or that $\partial_{\mu}Y=0$. The first of these implies
$H_{v}^{\prime\prime}(Y)=0\,\implies\,H_{v}(Y)=BY+A\,,$ (53)
which in the $\bar{\mu}=1$ frame, that is $H_{v}\rightarrow f_{v}/\sqrt{g}$ and $Y\rightarrow 1/\sqrt{g}$, reduces to the condition (18) on the coupling function $f_{v}$ we already found. The second option gives
$\partial_{\mu}Y=0\,\implies\,Y=\text{constant}\,,$ (54)
which in the $\bar{\mu}=1$ frame reduces to $g=$ constant, and this is precisely the condition (19) on the metric determinant we previously obtained.
Moreover, the EMT (48) in the potential regime takes the form
$T_{\mu\nu}=V\left(H_{v}-YH_{v}^{\prime}\right)g_{\mu\nu}\equiv\lambda\,g_{\mu\nu}\,,$ (55)
and if we use the above EoM we see that it is of a cosmological constant-type once again, since
$\lambda\equiv V\left(H_{v}-YH_{v}^{\prime}\right)=\text{constant}\,.$ (56)
Thus, in the potential limit we have recovered our previous results, as it should be expected.
IV.2 Kinetic domination in the covariantized approach
In the kinetic domination regime we neglect the potential terms. The EoM (46) for $T^{\mu}$ become
$\partial_{\mu}\left[H_{k}^{\prime}(Y)X\right]=\frac{1}{2}\partial_{\mu}\left[H_{k}^{\prime}(Y)(\partial\psi)^{2}\right]=0\,.$ (57)
It follows from the above that
$H_{k}^{\prime}(Y)(\partial\psi)^{2}=\text{constant}\equiv-c_{\rho}\,,$ (58)
where we have named the arbitrary constant as $-c_{\rho}$ for future convenience. On the other hand, the EoM (43) for $\psi$ in the kinetic domination regime becomes
$\nabla_{\mu}\left[H_{k}(Y)\partial^{\mu}\psi\right]=\nabla_{\mu}\left[H_{k}(Y)Nu^{\mu}\right]=0\,,$ (59)
where we have recalled the velocity (9). If we expand the equation and divide through by $H_{k}$ (which we take to be nonzero), then we obtain
$Nu^{\mu}\nabla_{\mu}\ln H_{k}+u^{\mu}\nabla_{\mu}N+N\nabla_{\mu}u^{\mu}=0\,.$ (60)
Recalling equation (23) for the relation between the expansion $\nabla_{\mu}u^{\mu}$ and the cross-sectional volume $\delta V$, and using the fact that the covariant derivatives may be changed by partial derivatives when acting on scalar functions, it follows from the above that
$\begin{split}Nu^{\mu}\partial_{\mu}\left(\ln H_{k}\right.&+\left.\ln N+\ln\delta V\right)=\\[5.0pt] &=Nu^{\mu}\partial_{\mu}\ln\left(H_{k}N\delta V\right)=0\,.\end{split}$ (61)
From this expression, we finally conclude that
$(\partial\psi)^{2}=\frac{C_{\psi}(x)}{(H_{k}\delta V)^{2}}\,,$ (62)
with $C_{\psi}(x)$ a function such that $u^{\mu}\partial_{\mu}C_{\psi}(x)=0$. Going now to the $\bar{\mu}=1$ frame and using equation (39) for $H_{k}|_{\bar{\mu}=1}$, we immediately recover expression (22) for the solution to the EoM in the TDiff approach.
On the other hand, substituting equation (62) into (58), it follows that
$-c_{\rho}=H_{k}^{\prime}(\partial\psi)^{2}=H_{k}^{\prime}\frac{C_{\psi}(x)}{(H_{k}\delta V)^{2}}\,,$ (63)
which leads to
$-\frac{H_{k}^{\prime}}{H_{k}^{2}}=\frac{c_{\rho}}{C_{\psi}(x)}\,\delta V^{2}\,.$ (64)
If we now define the function
$C_{g}(x)\equiv\frac{c_{\rho}}{C_{\psi}(x)}\,,$ (65)
which satisfies $u^{\mu}\partial_{\mu}C_{g}(x)=0$, we may write the above equation as
$-\frac{H_{k}^{\prime}}{H_{k}^{2}}=C_{g}(x)\delta V^{2}\,.$ (66)
It is now immediate to see that we recover the constraint (28) when going to the frame $\bar{\mu}=1$, since
$\left.-\frac{H_{k}^{\prime}}{H_{k}^{2}}\right|_{\bar{\mu}=1}=-\frac{f_{k}-2gf_{k}^{\prime}}{f_{k}^{2}\,/\,g}\,,$ (67)
and this implies that
$(2F-1)\frac{g}{f_{k}}=C_{g}(x)\delta V^{2}\,,$ (68)
where we have recalled the definition of the function $F$ in equation (27).
On another note, in the kinetic domination regime the EMT (48) takes the form
$T_{\mu\nu}=H_{k}\partial_{\mu}\psi\partial_{\nu}\psi-\left[H_{k}-YH_{k}^{\prime}\right]Xg_{\mu\nu}\,,$ (69)
which, under the assumption of the field derivative $\partial_{\mu}\psi$ being a timelike vector, is seen to be equivalent to that of a perfect fluid, using the same definition (9) for the velocity, and with energy density
$\rho=-\frac{c_{\rho}}{2}\left(\frac{H_{k}}{H_{k}^{\prime}}+Y\right)$ (70)
and pressure
$p=-\frac{c_{\rho}}{2}\left(\frac{H_{k}}{H_{k}^{\prime}}-Y\right)\,,$ (71)
where we have made use of the solution (58) to simplify $H_{k}^{\prime}X=-c_{\rho}/2$. Note that we have obtained that the energy density and pressure are functions of only $Y$, and this dependence on a single variable means that possible perturbations of the fluid will be adiabatic.
The EoS now reads
$w=\frac{p}{\rho}=\frac{\frac{H_{k}}{H_{k}^{\prime}}-Y}{\frac{H_{k}}{H_{k}^{\prime}}+Y}\,,$ (72)
and, as expected, is seen to be a function of $Y$ only. Note also that in the $\bar{\mu}=1$ frame we have
$\left.w\right|_{\bar{\mu}=1}=\frac{\frac{f_{k}/\sqrt{g}}{f_{k}-2gf^{\prime}_{k}}-\frac{1}{\sqrt{g}}}{\frac{f_{k}/\sqrt{g}}{f_{k}-2gf^{\prime}_{k}}+\frac{1}{\sqrt{g}}}=\frac{gf^{\prime}_{k}}{f_{k}-gf^{\prime}_{k}}\,,$ (73)
which coincides with expression (26) for the EoS previously found in the TDiff formalism. Consider now the following quantity:
$w-1=\frac{-2Y}{\frac{H_{k}}{H_{k}^{\prime}}+Y}=\frac{Yc_{\rho}}{\rho}\,,$ (74)
where in the second equality we have recalled expression (70) for the energy density. Rearranging, we obtain the following simple expression for the energy density:
$\rho=\frac{Yc_{\rho}}{w-1}\,.$ (75)
After evaluation in the frame $\bar{\mu}=1$, i.e. substituting $Y\rightarrow 1/\sqrt{g}$ and recalling from equation (73) that the EoS is recovered, it finally yields the simple expression (31) for the energy density that was obtained in the TDiff approach.
Finally, the speed of sound of the adiabatic perturbations follows from the definition $\delta p=c_{s}^{2}\,\delta\rho$, which joined with $p=w\rho$ yields
$c_{s}^{2}=w+w^{\prime}\frac{\rho}{\rho^{\prime}}\,.$ (76)
Substituting, it follows that
$c_{s}^{2}=\frac{H_{k}\,H_{k}^{\prime\prime}}{H_{k}\,H_{k}^{\prime\prime}-2\left(H_{k}^{\prime}\right)^{2}}\,.$ (77)
It is possible once again to verify that this expression, when evaluated in the $\bar{\mu}=1$ frame, reduces to the previously obtained speed of sound in equation (32). To this end, we must simply recall equations (39) and (50) for $H|_{\bar{\mu}=1}$ and $H^{\prime}|_{\bar{\mu}=1}$, respectively, and also use that
$H^{\prime\prime}(Y)=2Y^{-3}\left[\frac{df}{d\left(Y^{-2}\right)}+2Y^{-2}\frac{d^{2}f}{d^{2}\left(Y^{-2}\right)}\right]\,,$ (78)
which yields
$H^{\prime\prime}(Y)\bigg{|}_{\bar{\mu}=1}=2g^{3/2}\big{[}f^{\prime}(g)+2gf^{\prime\prime}(g)\big{]}\,.$ (79)
After careful substitution, we finally obtain the speed of sound in equation (32), that is
$\left.c_{s}^{2}\right|_{\bar{\mu}=1}=-\frac{gf_{k}(f^{\prime}_{k}+2gf_{k}^{\prime\prime})}{f_{k}^{2}+(2gf_{k}^{\prime})^{2}-gf_{k}(5f_{k}^{\prime}+2gf_{k}^{\prime\prime})}\,.$ (80)
We thus conclude that, proceeding via the covariantized approach, we have reobtained all of the results from the TDiff approach.
IV.3 Simple models in the kinetic regime
Regarding some particular models of interest in the kinetic regime, note that one can obtain a constant equation of state parameter for the fluid considering a power-law kinetic function. Indeed, taking into account equation (72) one has
$H_{k}(Y)=CY^{\beta}\,\implies\,w=\frac{1-\beta}{1+\beta}\,,$ (81)
As it was already discussed in reference [12] following the TDiff approach, in this case we have that the propagation speed of the field perturbations is
$c_{s}^{2}=\frac{1-\beta}{1+\beta}=w\,,$ (82)
which has been obtained taking into account equation (77). This is because a power-law kinetic function in the covariantized approach implies a power-law kinetic coupling in the TDiff approach but with a different value of the exponent (note equation (37)).
However, the models with a constant equation of state parameter are the only family that will have a similar functional expression in both approaches. For example, an exponential kinetic function in the covariantized approach is not equivalent to an exponential kinetic coupling in the TDiff approach according to equation (37). In this case, equation (72) leads to the equation of state parameter
$H_{k}(Y)=Ce^{\beta Y}\,\implies\,w=\frac{1-\beta Y}{1+\beta Y}\,,$ (83)
which could interpolate between the behaviour of stiff matter and that of a cosmological constant. Nevertheless, taking into account equation (77), one obtains
$c_{s}^{2}=-1\,.$ (84)
So, the particular models with an exponential kinetic function in the covariantized approach are unstable, whereas models with an exponential kinetic coupling in the TDiff approach can have interesting phenomenology [13]. Finally, thus, it should be emphasized that models which could be natural to consider in one approach (due to a simple form of the relevant function) could appear unnatural in the other approach (in which the relevant function could be more complicated).
V Conclusions
In this work we have explored the idea of restoring the Diff invariance to a theory in which it was previously broken down to TDiff. Inspired by the work of Stueckelberg [15] for gauge fields and the treatment by Henneaux and Teitelboim [17] for Unimodular Gravity, we have reformulated a TDiff invariant theory for a scalar field $\psi$ in a way in which the invariance of the theory under general diffeomorphisms is recovered at the expense of introducing an additional vector field $T^{\mu}$. We have highlighted that the two approaches are indeed equivalent ways of dealing with the same problem.
In the TDiff approach, the main tool was the study of the EMT conservation. Indeed, due to the lack of Diff invariance, it was not an automatic consequence of the symmetry, but rather a consistency condition which should be satisfied by the solutions of the theory. On the other hand, in the covariantized approach the symmetry is restored and the EMT is once again trivially conserved. In this case, then, the additional information is provided by the EoM of the newly incorporated vector field $T^{\mu}$. It turns out that the analysis of these EoM is rather straightforward, and allows one to reach the same results as in the TDiff approach.
The model we have considered in this work is a simple one; nevertheless, it has shown the great utility of the covariantized approach, and that is the main point of the paper. It is intriguing to consider what other problems and analyses, perhaps difficult in one approach, might the other approach help simplify. Moreover, having two separate yet equivalent ways of tackling the same problem provides one with a consistency check of sorts, in the sense that a result obtained via one approach may be compared with that obtained through the other.
Overall, we conclude that the covariantized treatment carried out in this work is a useful tool for several reasons. From the fundamental point of view, the restoration of a symmetry via the introduction of additional quantities can be useful in what respects discerning which are the truly dynamical components of a theory. This shall all be further pursued in future work. From the more practical point of view, we have seen how the more natural models in one approach are not necessarily those that one would consider in the other approach in the first place. So, one could be missing interesting phenomenology by focusing only in one of the approaches. For all of these reasons, we think that combining both approaches will allow us to better understand different aspects of the theory.
Acknowledgements.
The authors would like to thank an anonymous referee of reference [5] for useful comments and suggestions which sparked the analysis in this paper. This work has been supported by the MICIN (Spain) project PID2022-138263NB-I00 (AEI/FEDER, UE). DJG also acknowledges financial support from the Ayudas de Doctorado IPARCOS-UCM/2023.
Appendix A Covariantized EMT
We include in this Appendix a couple of calculations regarding the covariantized EMT which are not needed to follow the main thread of the discussion, but which could come in handy in order to clear up some of the computations.
The first one has to do with obtaining expression (48) for the covariantized EMT. As we know from definition (6), we must find the functional derivative of the matter action
$S_{m}=\int d^{4}x\,\sqrt{g}\left[H_{k}(Y)X-H_{v}(Y)V(\psi)\right]$ (85)
with respect to the metric. To this end, we shall find the variation $\delta S_{m}$ with respect to the metric, which reads
$\begin{split}\delta S_{m}&=\int d^{4}x\,\bigg{\{}\left(H_{k}X-H_{v}V\right)\delta\sqrt{g}+\sqrt{g}\,X\delta H_{k}\\ &\qquad+\sqrt{g}\,H_{k}\delta X-\sqrt{g}\,V\delta H_{v}-\sqrt{g}\,H_{v}\delta V\bigg{\}}\,.\end{split}$ (86)
The needed variations with respect to the metric are the following:
$\delta V$ $=0\,,$ (87a)
$\delta X$ $=\frac{1}{2}\partial_{\mu}\psi\partial_{\nu}\psi\,\delta g^{\mu\nu}\,,$ (87b)
$\delta\sqrt{g}$ $=-\frac{1}{2}\sqrt{g}\,g_{\mu\nu}\,\delta g^{\mu\nu}\,,$ (87c)
$\delta H$ $=H^{\prime}\left[\frac{1}{2}Yg_{\mu\nu}\,\delta g^{\mu\nu}+\frac{1}{\sqrt{g}}\partial_{\alpha}\left(T^{\alpha}\,\delta\sqrt{g}\right)\right]\,.$ (87d)
The usage of the first three is simple enough, but the fourth one deserves a bit more care. After substitution of everything into the integral (86) we will have a term of the form
$\begin{split}\sqrt{g}\,X\delta H_{k}&=\sqrt{g}\,XH_{k}^{\prime}\frac{1}{2}Yg_{\mu\nu}\,\delta g^{\mu\nu}-T^{\alpha}\delta\sqrt{g}\,\partial_{\alpha}\left(H_{k}^{\prime}X\right)\\ &\quad+\partial_{\alpha}\left(T^{\alpha}H_{k}^{\prime}X\delta\sqrt{g}\right)\end{split}$ (88)
and a term of the form
$\begin{split}\sqrt{g}\,V\delta H_{v}&=\sqrt{g}\,VH_{v}^{\prime}\frac{1}{2}Yg_{\mu\nu}\,\delta g^{\mu\nu}-T^{\alpha}\delta\sqrt{g}\,\partial_{\alpha}\left(H_{v}^{\prime}V\right)\\ &\quad+\partial_{\alpha}\left(T^{\alpha}H_{v}^{\prime}V\delta\sqrt{g}\right)\,.\end{split}$ (89)
A couple of steps are now needed for simplification. The first step is to, as usual, discard the boundary terms appearing in the last lines of the above two expressions. The second step is to recognize that after the combination $\sqrt{g}\,X\delta H_{k}-\sqrt{g}\,V\delta H_{v}$ we will obtain (among others) a term of the form
$T^{\alpha}\delta\sqrt{g}\,\partial_{\alpha}\left(H_{v}^{\prime}V-H_{k}^{\prime}X\right)=0\,,$ (90)
which vanishes by virtue of the EoM (46) for the vector field $T^{\mu}$. Carefully substituting and taking everything into account, we obtain the EMT (48) in the end, i.e.
$\begin{split}T_{\mu\nu}&=H_{k}(Y)\partial_{\mu}\psi\partial_{\nu}\psi-\left[H_{k}(Y)X-H_{v}(Y)V\right]g_{\mu\nu}\\[5.0pt] &\qquad+Y\left[H_{k}^{\prime}(Y)X-H_{v}^{\prime}(Y)V\right]g_{\mu\nu}\,.\end{split}$ (91)
The second comment we wish to make is regarding the conservation of this EMT, which should be trivial given that the theory has Diff symmetry, and it can be explicitly checked that this is so. Proceeding piece by piece, we have that
$\begin{split}\nabla_{\mu}\left(H_{k}\nabla^{\mu}\psi\nabla^{\nu}\psi\right)&=\nabla^{\nu}\psi\nabla_{\mu}\left(H_{k}\nabla^{\mu}\psi\right)\\[5.0pt] &\quad+H_{k}\nabla^{\mu}\psi\left(\nabla_{\mu}\nabla^{\nu}\psi\right)\,,\end{split}$ (92a)
$\begin{split}\nabla_{\mu}\left[\left(H_{k}X-H_{v}V\right)g^{\mu\nu}\right]&=H_{k}\nabla^{\mu}\psi\left(\nabla_{\mu}\nabla^{\nu}\psi\right)\\[5.0pt] &\quad+\nabla^{\nu}Y\left(H_{k}^{\prime}X-H_{v}^{\prime}V\right)\\[5.0pt] &\quad-H_{v}V^{\prime}\nabla^{\nu}\psi\,,\end{split}$ (92b)
$\begin{split}\nabla_{\mu}\left[Y\left(H_{k}^{\prime}X-H_{v}^{\prime}V\right)g^{\mu\nu}\right]&=\nabla^{\nu}Y\left(H_{k}^{\prime}X-H_{v}^{\prime}V\right)\\[5.0pt] &\quad+Y\nabla^{\nu}\left(H_{k}^{\prime}X-H_{v}^{\prime}V\right)\,,\end{split}$ (92c)
Joining now all the pieces together we can see some very convenient cancellations, and also that
[…]
i.e. we indeed obtain that the conservation of the EMT is trivially satisfied on the solutions to the EoM of the theory, as we expected.